OpenStudy (anonymous):

WILL GIVE MEDAL TO BEST ANSWER How does finding the square root of a number compare to finding the cube root of a number? Use the number 64 in your explanation.

OpenStudy (jdoe0001):

\(\large {a\cdot a=x\qquad then\qquad \sqrt{x}=a\\ \quad \\ a\cdot a\cdot a=y\qquad then\qquad \sqrt[3]{y}=a}\)